# Between two descriptions of dark matter around a black hole: photon sphere, shadow, and lensing

### v1

@@ Abstract
We examine observational discrepancies of two widely used models describing anisotropic (dark) matter distributions around a black hole, focusing on their photon spheres, shadow radii, and lensing observables. The models considered are the vacuum and Einstein cluster dark matter model, characterized by negative and zero radial pressure, respectively. The analysis reveals that these models display contrasting photon sphere behavior. In particular, the Einstein cluster results in a more pronounced deviation in the shadow radius relative to the standard Schwarzschild black hole. Additionally, a distinctive lensing phenomenon associated with the matter halo is identified in both models.
@@ I Introduction
Black holes (BHs) at galactic centers are expected to be surrounded by dark matter (DM) distributions [1, 2]. A wide range of astrophysical observations provides strong evidence for the presence of DM in galaxies (see, e.g., [3, 4]). Motivated by this, numerous theoretical studies have explored the influence of DM-surrounded BHs, including the observational signatures from such systems [5, 6, 9, 12, 11, 13, 10, 19, 7, 8, 18, 14, 15, 17, 16, 20, 21, 22, 23]. These studies typically use phenomenological approaches to model DM distributions, some of which are constructed based on N-body simulations of galaxies [24].
However, there is still no consensus on how DM modifies the spacetime geometry in the vicinity of a BH. A number of studies adopt the simplifying assumption of a vacuum spacetime [13, 12, 14, 18, 17, 15, 16],
@@
effectively modeling the region around the BH as being oblivious to the presence of DM. It has been argued that this approximation captures the essential physical behavior of more general treatments [5]. In contrast, Cardoso et al. [6] used the Einstein cluster to model the (dark) matter distributions around a BH in a galaxy, which leads to a spacetime geometry that explicitly violates condition (1) due to the zero pressure matter distribution. This distinction raises the question: Does the simplified metric approximation adequately capture the observational signatures expected from a more complete Einstein cluster description of DM in the strong field regime?
This letter addresses the above question from the perspective of electromagnetic observations. While other studies focus on a single DM description, our work compares the two models and identifies their key distinguishing features. We consider two observational regimes. First, we examine the BH shadow as a probe of the strong-field region. Although numerous studies have examined this context, most of them employ the assumption (1[9, 12, 11, 13, 14, 10]. This approach may provide insight into the characteristics of galactic DM inferred from the shadow of the central BH, particularly given observational results from the Event Horizon Telescope [25, 26]. Second, motivated by the fact that DM distributions typically span over galactic scales, we investigate a complementary observable in a weaker-field regime—namely, the image positions and time delays arising from gravitational lensing, following the formalism of Virbhadra [27]. This choice was motivated by the findings of Boehmer and Harko [28] who reported that the Newtonian approximation of the DM-induced deflection angle is about $\sim 77\%$ weaker than predictions from the Einstein cluster description of DM in the constant velocity region of galaxies. Kouniatalis et al. [19] recently employed the Einstein cluster to investigate its impact on light deflection near the BH. In contrast, our work focuses on directly measurable lensing observables, as outlined above.
@@
When modeling the DM distribution as an anisotropic fluid with $T_{\mu}^{\nu}=\text{diag}(-\epsilon,p,p_{t},p_{t})$, there is no general consensus on the appropriate equation of state (EoS) for the radial pressure $p$. A number of studies adopt, for simplicity, the implicit choice $p=-\epsilon$, leading to the energy–momentum tensor
@@
We refer to this particular EoS as vacuum DM, since it is equivalent to a de Sitter–type EoS. This choice is, in fact, a consequence of the ansatz (1). It has been argued that this condition yields similar physical effects to more general cases where $g_{tt}\neq g_{rr}^{-1}$ in the presence of DM [5]. Moreover, this ansatz provides a straightforward, widely used procedure for constructing rotating BH solutions via the Newman–Janis algorithm applied to the seed metric [17, 15, 16].
@@
This DM model is known as the Einstein cluster (EC), which describes a self-gravitating matter distribution composed of particles on stable circular orbits around the center [28], as originally proposed by Einstein [29]. The system is stabilized by the balance between gravity and the centrifugal forces from the orbiting matter, resulting in an anisotropic energy–momentum tensor with zero radial pressure. Unlike the vacuum DM case, a rotating solution for EC DM cannot be obtained via the Newman–Janis algorithm; its numerical rotating solution appeared only recently in Ref. [30]. Moreover, a recent study suggests that EC DM arises as a solution to a non-minimally coupled vector field interacting with gravity [31], indicating it may effectively behave as a vector field.
@@
We now turn to the profile shape of the energy density $\epsilon(r)$, where most studies adopt effective distribution profiles. In this work, we employ the general form as in Ref. [8]
@@
Earlier studies have reported the vanishing DM energy density near the BH horizon [1, 2]. Analytic approximations to this effect have been constructed by incorporating a cut-off factor into the energy density profile. For instance, the authors of Refs. [6, 7, 8] use
@@
where $r_{c}$ denotes the cut-off radius at which the DM density vanishes. We refer to Eq. (6) as the modified energy density profile. In Ref. [6], the authors assumed $r_{c}$ to coincide with the BH horizon ($r_{c}=2M_{BH}$). However, this assumption may lead to violations of the energy conditions and of the transverse speed of sound [34], motivating Ref. [7] to adopt a larger value of $r_{c}$—specifically $r_{c}=4M_{BH}$, following the results of Ref. [2].
@@ III Spacetime geometry
$m_{DM}(r)=\int_{r_{c}}^{r}4\pi\epsilon(\tilde{r}),d\tilde{r}$. (8)
The total Arnowitt–Deser–Misner (ADM) mass of the system is then $M_{ADM}=M_{BH}+M_{DM}$, where $M_{DM}=\lim_{r\to\infty}m_{DM}(r)$. The full analytic expressions of the mass functions corresponding to the modified Hernquist and Dehnen–$3/2$ profiles are provided in the Supplementary Material.
Throughout our analysis, we consider two representative values for the DM total mass and core size: $M_{DM}/M_{BH}\in\{10,100\}$ and $a_{0}/M_{BH}\in\{10^{3},10^{4}\}$. These choices are sufficient for our purposes, as they represent both diluted and relatively compact DM halos. We also adopt the same cut-off radius $r_{c}=2M_{BH}$ as in Ref. [6].
Figure 1: The shift function for the EC DM with $M_{DM}=100M_{BH}$. The gray dashed line represents the vacuum DM case, for which $\Phi(r)=0$ everywhere.
Now we are left with the time component, in particular the shift function $\Phi(r)$. Following the formalism in Ref. [35], the shift function for each DM case satisfies
@@
Figure 2: (Top) Photon sphere radius and (Bottom) shadow radius for the EC DM and vacuum DM as functions of the DM mass, shown by solid and dotted lines, respectively.
Figure 3: (Top) Separation angle between the lensed primary and secondary images as a function of $\alpha$, and (Bottom) their differential time delay, for EC DM (solid lines) and vacuum DM (dotted lines).
Figure 4: (Top) Angular positions of the primary and secondary images as functions of $\alpha$, and (Bottom) their corresponding time delays, for EC DM (solid lines) and vacuum DM (dotted lines), with $M_{DM}/M_{BH}=100$ and $D_{LS}/a_{0}=1000$.
We show the non-zero shift function for the EC DM in Fig. 1, where we set $M_{DM}/M_{BH}=100$ for better visualization. The plot illustrates that the additional redshift factor extends beyond $a_{0}$, and that a more compact configuration (smaller $a_{0}$) leads to a larger deviation from the zero shift function over a shorter radial range. Moreover, the Dehnen 3/2 profile produces a substantially stronger redshift factor near the BH horizon compared to the Hernquist profile, indicating that the discrepancy between EC DM and vacuum DM is likely model dependent. As we will show, this additional redshift factor plays a crucial role in electromagnetic observations, and its impact becomes non-negligible in the strong-field regime.
@@ IV Resulting observables
Since we are dealing with a static and spherically symmetric spacetime, the geodesic equations for photons take a simple form. The two relevant equations are
$\frac{d\phi}{dr}=$ $\pm\frac{1}{r^{2}}\left[\frac{-g_{tt}g_{rr}}{b^{-2}-V(r)}\right]^{1/2},$ (10)
$\frac{dt}{dr}=$ $\pm\frac{1}{b}\left[\frac{g_{rr}(-g_{tt})^{-1}}{b^{-2}-V(r)}\right]^{1/2},$ (11)
where $V(r)\equiv-g_{tt}/r^{2}$ is the photon effective potential, $b\equiv L/E$ is the impact parameter, $E$ and $L$ denoting the photon’s energy and angular momentum, respectively. The effective potential determines the turning point of photons approaching the system, and its form depends explicitly on $g_{tt}$. Therefore, we expect EC DM to yield electromagnetic observables that differ from those obtained in the vacuum DM approximation.
Photon sphere and shadow—To calculate the BH shadow radius, one first needs to determine the location of the photon sphere $r_{ps}$. This is obtained by extremizing the effective potential, i.e., by solving $\partial_{r}V|_{r=r_{ps}}=0$. Once the photon sphere is located, the critical impact parameter $b_{c}$ follows from the relation $b_{c}=V(r_{ps})^{-1/2}$. Since the spacetime is asymptotically flat, the shadow radius $R_{sh}$, as well as its angular radius $\Theta_{sh}$ as seen by a distant observer at $D_{OL}\to\infty$, are given by [37]
$R_{sh}\approx b_{c},\qquad\Theta_{sh}\approx\frac{R_{sh}}{D_{OL}}.$ (12)
The resulting photon sphere radius, together with the BH shadow radius, is shown in Fig. 2. One can clearly observe that the photon sphere exhibits opposite trends in the two DM models: increasing the total DM mass shifts the photon sphere outward in the EC DM case, whereas in the vacuum DM case, it is shifted inward. The latter behavior commonly appears in models that adopt a de Sitter–like vacuum equation of state to describe the effective matter sourcing the geometry, as is typical in regular BH solutions (see, e.g., Refs . [38, 39]). Nonetheless, the deviation of the photon sphere from its Schwarzschild value ($r_{ps}=3M_{BH}$) remains extremely small, even in the most compact configurations with small $a_{0}$ and large $M_{DM}$.
@@
This point is crucial when using BH shadow observations to constrain DM parameters. For example, by simply applying the M87* shadow radius constraint from Refs. [40, 41] to the Hernquist distribution with $a_{0}/M_{BH}=10^{3}$, the EC DM model sets an upper bound on total DM mass at $M_{DM}/M_{BH}\approx 150$. In contrast, the vacuum DM model permits a much greater DM mass, up to the threshold where the halo forms a secondary horizon. This is consistent with Ref. [12] who found that notable changes in the BH shadow radius in vacuum DM models appear only when DM is highly concentrated near the BH, i.e., when $a_{0}\sim\sqrt{3M_{BH}M_{DM}}$. Thus, the DM description—whether as an Einstein cluster or a vacuum fluid—must be carefully chosen, as it yields qualitatively different astrophysical conclusions.
@@
Lensing observables—Next, we investigate the discrepancy between the two DM models in their lensing observables. Similar analyses of DM effects on lensing have recently been carried out in Ref. [19] for the EC DM model with the modified Hernquist profile of Ref. [6]. Numerous studies have also explored lensing signatures in various vacuum DM scenarios (see, for example, Refs. [20, 21, 22, 23] for some of the latest results).
@@
The gravitational lensing setup is described as follow. A source located at an angular position $\alpha$ relative to the observer produces two observable images located at an angle $\Theta^{(+)}$ and $\Theta^{(-)}$, corresponding to the primary and secondary images, respectively [27]. For each image, the difference between the lapsed travel time in the curved spacetime and that in flat spacetime defines the time delay $t_{d}^{(\pm)}$. Since the primary $(+)$ and secondary ()(-) rays follow different paths, they yield distinct time delays. The directly measurable quantities are the angular separation between the primary and secondary images, $\Delta\Theta=|\Theta^{(+)}-\Theta^{(-)}|$, and their differential time delay, $\Delta t_{d}=|t_{d}^{(+)}-t_{d}^{(-)}|$. An illustration and detailed description of the computation procedure of these observables are provided in the Supplementary Material.
@@
For our purposes, we assume the source lies within or near the DM halo at a distance $D_{LS}$ from the BH, while the observer is at $D_{OL}$ from the BH and positioned in a region where spacetime is effectively flat ($g_{rr}\approx-g_{tt}\approx 1$). This setup ensures a significant influence of the redshift factor, particularly on the primary images. If both $D_{OL}$ and $D_{LS}$ were very large, light rays of the primary images would approach closest to the BH outside the DM halo, where spacetime resembles a Schwarzschild geometry and the DM-induced redshift factor is negligible. We specifically set $D_{OL}/M_{BH}\approx 10^{11}$, and choose the source distance from the lens as $D_{LS}/a_{0}\in\{10,1000\}$. Observable quantities are presented in arbitrary units to facilitate the comparison between the two DM descriptions. As a reference, using Eq. (12), the shadow angular radius $\Theta_{sh}$ for a Schwarzschild BH in this setup is $\sim 10.7,\mu\text{as}$, while the angular radius of the DM core with $a_{0}/M_{BH}=10^{3}$ is $\sim 2.06\times 10^{3}\,\mu\text{as}$.
@@
We compare the EC and vacuum DM predictions for both observables in Fig. 3, focusing on the case with $M_{DM}/M_{BH}=10$. In all configurations, EC DM produces stronger lensing than vacuum DM for image separation and differential time delay. The largest differences on the image position appear when the source, lens, and BH are aligned ($\alpha=0$), resulting in the so-called Einstein ring [42]. The size of this discrepancy depends on both the source distance $D_{LS}$ and the DM core size. A smaller core causes more significant differences for sources near the BH, while a larger core amplifies the effect for more distant sources. EC DM predicts an Einstein ring diameter $\sim 4\%-56\%$ larger than vacuum DM, depending on configuration and source distance.
@@
In the setup with $M_{DM}/M_{BH}=100$ and $D_{LS}/a_{0}=10^{3}$, in Fig. 4, we find an interesting phenomenon arising from the DM distribution: there exists a range of $\alpha$ that produces multiple secondary images across all DM models. This phenomenon was also reported in a recent lensing study of EC DM around a BH [19] on the deflection angle. A similar feature is clearly visible in Fig. 4c of Ref. [11], although they consider a different DM and lensing configuration. The discrepancy between the two DM descriptions is reflected in the positions and separations of the multiple secondary images. Vacuum DM produces multiple lensed images at smaller source inclinations from the observer, but with a larger separation than those of the EC DM. However, the time delays associated with these multiple lensed images are roughly the same, implying that all events in these lensed images occurred at the same time as observed by a distant observer. This may become an important potential signature of DM around a BH, regardless of the chosen description of the DM halo itself.

### v4

# Two descriptions of dark matter around a black hole: photon sphere, shadow, and lensing

@@ Abstract
We examine the observational discrepancies of two widely used models describing anisotropic (dark) matter distributions around a black hole, focusing on their photon spheres, shadow radii, and lensing observables. The models considered are the vacuum and Einstein cluster dark matter models, characterized by negative and zero radial pressure, respectively. The analysis reveals that these models display contrasting photon sphere behaviors. In particular, the Einstein cluster results in a more pronounced deviation in the shadow radius relative to the standard Schwarzschild black hole. Additionally, a distinctive lensing phenomenon associated with the matter halo is identified in both models.
@@ I Introduction
Black holes (BHs) at galactic centers are expected to be surrounded by dark matter (DM) distributions [1, 2, 3, 4]. A wide range of astrophysical observations provides strong evidence for the presence of DM in galaxies (see, e.g., [6, 5]). Motivated by this, numerous theoretical studies have explored the influence of DM-surrounded BHs, including the observational signatures from such systems [11, 10, 17, 25, 8, 16, 7, 9, 12, 24, 27, 18, 28, 26, 34, 13, 14, 15, 33, 29, 30, 32, 31, 20, 21, 22, 23, 19]. These studies typically use phenomenological approaches to model DM distributions —i.e., using empirical DM models (e.g. [35, 36, 1, 37]) instead of solutions derived from fundamental principles—some of which were constructed based on N-body simulations of galaxies.
However, there is still no consensus on how DM modifies the spacetime geometry in the vicinity of a BH. A number of studies adopt the simplifying assumption of a vacuum spacetime [10, 11, 9, 8, 28, 27, 29, 20, 21, 22, 23, 33, 32, 30, 31, 24, 25, 26],
@@
effectively modeling the region around the BH as being oblivious to the presence of DM. It has been argued that this approximation captures the essential physical behavior of more general treatments [9]. This approach was recently criticized in Ref. [38], as the metric is inconsistent with the intended matter distribution. In contrast, Cardoso et al. [12] used the Einstein cluster to model the (dark) matter distributions around a BH in a galaxy, which leads to a spacetime geometry that explicitly violates condition (1) due to the zero radial pressure matter distribution. This distinction raises the question: Does the simplified metric approximation adequately capture the observational signatures expected from a more complete Einstein cluster description of DM in the strong field regime?
This letter addresses the above question from the perspective of electromagnetic observations. While other studies focus on a single DM description, our work compares the two models and identifies their key distinguishing features. We consider two observational regimes. First, we examine the BH shadow as a probe of the strong-field region. Although numerous studies have examined this context [24, 27, 28, 29, 26, 25, 19, 18, 17, 16], most of them employ the assumption (1)—or an equivalent condition leading to (1[24, 27, 28, 29, 26, 25]. This approach may provide insight into the characteristics of galactic DM inferred from the shadow of the central BH, particularly given observational results from the Event Horizon Telescope [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Second, motivated by the fact that DM distributions typically span over galactic scales, we investigate a complementary observable in a weaker-field regime—namely, the image positions and time delays arising from gravitational lensing, following the formalism of Virbhadra [56]. This choice was motivated by the findings of Boehmer and Harko [57] who reported that the Newtonian approximation of the DM-induced deflection angle is about $\sim 77\%$ weaker than predictions from the Einstein cluster description of DM in the constant velocity region of galaxies. Kouniatalis et al. [34] recently employed the Einstein cluster to investigate its impact on light deflection near the BH; in contrast, our work focuses on directly measurable lensing quantities, as outlined above, as well as comparing two different DM descriptions and energy density distributions profiles to assess their impact on the resulting spacetime geometry.
@@
When modeling the DM distribution as an anisotropic fluid with $T_{\mu}^{\nu}=\text{diag}(-\epsilon,p,p_{t},p_{t})$, there is no general consensus on the appropriate equation of state (EoS) for the radial pressure $p$. A number of studies adopt, for simplicity, the implicit choice $p=-\epsilon$ [10, 11, 9, 8, 28, 27, 29, 33, 32, 30, 31, 24, 25, 26], leading to the energy–momentum tensor
@@
We refer to this particular EoS as vacuum DM, since it is equivalent to a de Sitter–type EoS. This choice is, in fact, a consequence of the ansatz (1). It has been argued that this condition yields similar physical effects to more general cases where $g_{tt}\neq g_{rr}^{-1}$ in the presence of DM [9]. Moreover, this ansatz provides a straightforward, widely used procedure for constructing rotating BH solutions via the Newman–Janis algorithm [58] applied to the seed metric [9, 29, 32, 30, 31]; however, this algorithm was criticized in Ref. [59], as it introduces pathologies and inconsistencies into the final solutions (see also Refs. [60, 61, 62, 63]).
@@
This DM model is known as the Einstein cluster (EC), which describes a self-gravitating matter distribution composed of particles on stable circular orbits around the center [57], as originally proposed by Einstein [64]. The system is stabilized by the balance between gravity and the centrifugal forces from the orbiting matter, resulting in an anisotropic energy–momentum tensor with zero radial pressure. The (numerical) rotating BH immersed in an EC DM solution appeared only recently in Ref. [65]. Moreover, a recent study suggests that EC DM arises as a solution to a non-minimally coupled vector field interacting with gravity [66], indicating it may effectively behave as a vector field.
@@
We now turn to the profile shape of the energy density $\epsilon(r)$, for which most studies adopt semi-analytic distribution profiles obtained from, e.g., Refs. [35, 36, 1, 37], that effectively explain galaxy rotation curves. In this work, we employ the general form as in Ref. [14]
@@
Earlier studies have reported the vanishing DM energy density near the BH horizon, both by Newtonian-based analysis [2] and general relativistic treatment [3]. Analytic approximations to this effect have been constructed by incorporating a cut-off factor into the energy density profile. For instance, the authors of Refs. [12, 13, 14] use
@@
where $r_{c}$ denotes the cut-off radius at which the DM density vanishes. We refer to Eq. (6) as the modified energy density profile. In Ref. [12], the authors assumed $r_{c}$ to coincide with the BH horizon ($r_{c}=2M_{BH}$). However, this assumption may lead to violations of the energy conditions and of the transverse speed of sound [67], motivating Ref. [13] to adopt a larger value of $r_{c}$—specifically $r_{c}=4M_{BH}$, following the relativistic treatment results of Ref. [3].
@@ III Spacetime geometry
$m_{DM}(r)=\int_{r_{c}}^{r}4\pi\tilde{r}^{2}\epsilon(\tilde{r})d\tilde{r}$. (8)
The total Arnowitt–Deser–Misner (ADM) mass of the system is then $M_{ADM}=M_{BH}+M_{DM}$, where $M_{DM}=\lim_{r\to\infty}m_{DM}(r)$. The derivation and full analytic expressions of the mass functions corresponding to the modified Hernquist and Dehnen–$3/2$ profiles are provided in the Supplementary Material.
Throughout our analysis, we consider two representative values for both the total DM mass and the core size, each expressed relative to the BH mass: $M_{DM}/M_{BH}\in\{10,100\}$ and $a_{0}/M_{BH}\in\{10^{3},10^{4}\}$. These choices are sufficient for our purposes, as they represent both relatively diluted and compact DM halos. The ratio $a_{0}/M_{DM}$ roughly characterizes the compactness of the DM distribution.
Although we mentioned that a general relativistic treatment indicates $r_{c}=4M_{BH}$, we instead adopt the same cut-off radius $r_{c}=2M_{BH}$ as in Ref. [12]. This choice ensures that the DM effect extends into the vicinity of the BH horizon, as we aim to identify the most significant discrepancies between the two DM descriptions in the strong-field limit.
Figure 1: The shift function for the EC DM with $M_{DM}=100M_{BH}$ for the Hernquist and Dehnen-3/2 energy density profile. The gray dashed line represents the vacuum DM case, for which $\Phi(r)=0$ everywhere. The Dehnen-3/2 distribution results in a substantially higher redshift than the Hernquist profile near the BH horizon. In general, for a generalized Dehnen profile, the shift function near the BH horizon decreases as $\gamma_{D}$ increases.
Now we are left with the time component, in particular the shift function $\Phi(r)$. Following the formalism in Ref. [68], the shift function for each DM case satisfies (see the Supplementary Material)
@@
We compare the non-zero shift function induced by the EC DM for both the Hernquist and Dehnen-3/2 profiles in Fig. 1, where we set $M_{DM}/M_{BH}=100$ for better visualization. We obtain the shift function by numerically integrating Eq. (9) backward from (numerical) infinity, imposing the aforementioned boundary condition for an asymptotically flat spacetime. The plot illustrates that the additional redshift factor induced by the EC DM remains nonzero for radii $r\gtrsim a_{0}$, although the main deviation from the zero shift function occurs predominantly at scales of the order of $a_{0}$ or smaller. A more compact configuration (smaller $a_{0}/M_{\rm BH}$) leads to a larger deviation from the zero shift function over a shorter radial range. Moreover, the Dehnen-3/2 profile produces a substantially stronger redshift factor near the BH horizon compared to the Hernquist profile, indicating that the discrepancy between EC DM and vacuum DM is likely model dependent. As we will show, this additional redshift factor plays a crucial role in electromagnetic observations, and its impact becomes non-negligible in the strong-field regime.
@@ IV Resulting observables
Since we are dealing with a static and spherically symmetric spacetime, the geodesic equations for photons take a simple form. We provide the expression of the geodesic equation and its derivation in the Supplementary Material. The corresponding photon effective potential takes the form
$V(r)=\frac{-g_{tt}}{r^{2}}=\left[1-\frac{2m(r)}{r}\right]\frac{e^{\Phi}}{r^{2}}.$ (10)
This potential determines the turning point of photons approaching the system. Since its form depends explicitly on the shift function, we expect the EC DM scenario to yield electromagnetic observables that differ from those obtained under the vacuum DM approximation.
Figure 2: (Top) Photon sphere radius and (Bottom) shadow radius for the EC DM and vacuum DM as functions of the DM mass, shown by solid and dotted lines, respectively. The photon sphere radius behaves in opposite ways in the EC and vacuum DM cases. The BH shadow radius is significantly more sensitive to the DM mass in the EC DM scenario than in the vacuum DM case; a closer look at the vacuum DM shows that the shadow radius barely deviates from the Schwarzschild BH value of $3\sqrt{3}M_{BH}\approx 5.196M_{BH}$.
Photon sphere and shadow—To calculate the BH shadow radius, one first needs to determine the location of the photon sphere $r_{ps}$. This is obtained by extremizing the effective potential, i.e., by solving $\partial_{r}V|_{r=r_{ps}}=0$. Since the $\Phi(r)$ itself is obtained numerically, we solve the $r_{ps}$ by a generic numerical root-finding method. Once the photon sphere is located, the critical impact parameter $b_{c}$ follows from the relation $b_{c}=V(r_{ps})^{-1/2}$. Since the spacetime is asymptotically flat, the shadow radius $R_{sh}$ as seen by a distant observer is given by [69]
$R_{sh}\approx b_{c}.$ (11)
The resulting photon sphere radius, together with the BH shadow radius, is shown in Fig. 2. One can clearly observe that the photon sphere exhibits opposite trends in the two DM models: increasing the total DM mass shifts the photon sphere outward in the EC DM case, whereas in the vacuum DM case, it is shifted inward. The latter behavior commonly appears in models that adopt a de Sitter–like vacuum equation of state to describe the effective matter sourcing the geometry, as is typical in regular BH solutions (see, e.g., Refs . [70, 71]). Nonetheless, the deviation of the photon sphere from its Schwarzschild value ($r_{ps}=3M_{BH}$) remains extremely small, even in the most compact configurations with small $a_{0}/M_{DM}$.
@@
This point is crucial when using BH shadow observations to constrain DM parameters. For example, by simply applying the $2\sigma$ Sagittarius A* shadow radius constraint from Ref. [72] (see also Ref. [53]),
$4.21M_{BH}\lesssim R_{sh}^{(\text{Sgr A*})}\lesssim 5.56M_{BH},$ (12)
@@
to the Hernquist distribution with $a_{0}/M_{BH}=10^{3}$, the EC DM model sets an upper bound on total DM mass at $M_{DM}/M_{BH}\approx 67$. In contrast, the vacuum DM model permits a much greater DM mass, up to the threshold where the halo forms a secondary horizon. This is consistent with Ref. [27] who found that notable changes in the BH shadow radius in vacuum DM models appear only when DM is highly concentrated near the BH, i.e., when $a_{0}\sim\sqrt{3M_{BH}M_{DM}}$. Thus, the DM description—whether as an Einstein cluster or a vacuum fluid—must be carefully chosen, as it yields qualitatively different astrophysical conclusions.
Figure 3: (Top) Separation angle between the lensed primary and secondary images as a function of $\alpha$, and (Bottom) their differential time delay, for EC DM (solid lines) and vacuum DM (dotted lines). As a reference, the shadow angular diameter of a Schwarzschild BH in this setup is $\sim 21.4\,\mu\text{as}$. Significant discrepancies in the separation angle between EC DM and vacuum DM are observed at the Einstein ring (when $\alpha=0$). On the other hand, the differential time delay shows the largest differences at higher source inclinations.
Figure 4: (Top) Angular positions of the primary and secondary images as functions of $\alpha$, and (Bottom) their corresponding time delays, for EC DM (solid lines) and vacuum DM (dotted lines), with $M_{DM}/M_{BH}=100$ and $D_{LS}/a_{0}=1000$. At certain source inclinations in both the EC DM and vacuum DM cases, the DM produces multiple secondary images. A closer examination of their corresponding time delays indicates that these images have similar time delays.
@@
Lensing observables—Next, we investigate the discrepancy between the two DM models in their lensing observables. Similar analyses of DM effects on lensing have recently been carried out in Ref. [34] for the EC DM model with the modified Hernquist profile of Ref. [12]. Numerous studies have also explored lensing signatures in various DM scenarios (see, for example, Refs. [19, 20, 21, 22, 23] for some of the latest results).
@@
The gravitational lensing setup is described as follows. A source at angular position $\alpha$ relative to the observer produces two images at angles $\Theta^{(+)}$ and $\Theta^{(-)}$, corresponding to the primary and secondary images, respectively [56]. The time delay for each image, $t_{d}^{(\pm)}$, is defined as the difference between the travel time in curved spacetime and that in flat spacetime. The observable quantities are the angular separation $\Delta\Theta=|\Theta^{(+)}-\Theta^{(-)}|$ and the differential time delay $\Delta t_{d}=|t_{d}^{(+)}-t_{d}^{(-)}|$. An illustration and computational details are provided in the Supplementary Material.
@@
For our purposes, we assume the source lies within or near the DM halo at a distance $D_{LS}$ from the BH, while the observer is located at $D_{OL}$ in a region where spacetime is effectively flat ($g_{rr}\approx-g_{tt}\approx 1$). This setup ensures a significant influence of the DM-induced redshift, particularly on the primary images. If both $D_{OL}$ and $D_{LS}$ were very large, the closest approach of the primary rays would occur outside the DM halo, where the spacetime is approximately Schwarzschild and the redshift effect becomes negligible. We set $D_{OL}/M_{BH}\approx 10^{11}$ and choose the $D_{LS}$ relative to the DM core size $a_{0}$ as $D_{LS}/a_{0}\in\{10,1000\}$.
For reference, while adopting the mass of Sagittarius A* ($\sim 4.3\times 10^{6}M_{\odot}$ [73]) for $M_{BH}$ in our settings, we obtain $D_{OL}\approx 20\,\text{kpc}$, $D_{LS}\in(0.002,2)\,\text{pc}$, and $a_{0}\approx\{2\times 10^{-4},2\times 10^{-3}\}\,\text{pc}$. Obviously, these choices are not intended to represent the realistic astrophysical environment of Sagittarius A*. We instead adopt this configuration to illustrate the qualitative behavior of both models and, to avoid confusion, use the mass of Sagittarius A* to fix the units of our results.
@@
We compare the EC and vacuum DM predictions for both observables in Fig. 3, focusing on the case with $M_{DM}/M_{BH}=10$. In all configurations, EC DM produces stronger lensing effect than vacuum DM for both image separation and differential time delay. The largest differences in the image position between the vacuum DM and EC DM cases appear when the source, lens, and BH are aligned ($\alpha=0$), resulting in the so-called Einstein ring [74]. The size of this discrepancy depends on both the source distance $D_{LS}$ and the DM core size. A smaller core causes more significant differences for sources near the BH, while a larger core amplifies the effect for more distant sources.
The smallest and largest differences in the angular position of the Einstein ring between the EC DM and vacuum DM cases in our comparison occur at $D_{LS}/a_{0}=1000$. The smallest difference appears for the more compact DM configuration described by the Dehnen-3/2 profile ($2.52\times 10^{4}\,\mu\text{as}$ for EC DM and $2.42\times 10^{4}\,\mu\text{as}$ for vacuum DM), while the largest occurs for the less compact DM configuration described by the Hernquist profile ($6.06\times 10^{4}\,\mu\text{as}$ for EC DM and $4.78\times 10^{4}\,\mu\text{as}$ for vacuum DM). Therefore, our results suggest that EC DM predicts an Einstein ring diameter that is roughly $\sim 4\%-27\%$ larger than that predicted by vacuum DM.
Our results also show that the discrepancy in the differential time delay between the vacuum and EC DM cases is insignificant given the current observational capabilities. We predict a discrepancy on a timescale of less than an hour; even the $\sim 2-3$ hour delay reported for the Einstein Cross Q2237+0305 remains unreliable, as it is associated with relatively large uncertainties [75, 76]. This indicates that we may not be able to distinguish between the vacuum and EC DM models based on the observed differential time delay alone.
In the setup with $M_{DM}/M_{BH}=100$ and $D_{LS}/a_{0}=10^{3}$, in Fig. 4, we find an interesting phenomenon arising from the DM distribution: there exists a range of $\alpha$ that produces multiple secondary images across all DM models. This phenomenon was also reported in a recent lensing study of EC DM around a BH [34] on the deflection angle. A similar feature is clearly visible in Fig. 4c of Ref. [18]; however, the multiplication of secondary images in our case occurs for a significantly less compact DM configuration, at angular scales several orders of magnitude larger than the shadow boundary. Our analysis, supplemented by the results of Ref. [34], shows that the occurrence of multiple secondary images is a consequence of a range of impact parameters that yield roughly the same deflection angle (see the Supplementary Material).
@@
The discrepancy in multiple secondary images between the two DM descriptions is reflected in their positions and separations. The vacuum DM case produces multiple lensed images at smaller source inclinations as seen by the observer, but with an overall larger separation than in the EC DM case: for the Hernquist profile DM, the vacuum DM and EC DM scenarios yield the largest separation of the multiple secondary images of approximately $\sim 0.45\times 10^{5}\,\mu\text{as}$ and $\sim 0.22\times 10^{5}\,\mu\text{as}$, respectively (see the top inset of Fig. 4). However, the time delays associated with these multiple lensed images are roughly the same, implying that all events in these lensed images occurred at the same time as observed by a distant observer. This may become an important potential signature of compact DMs around a BH, regardless of the chosen description of the DM halo itself.